# Duality theory in linear optimization and its extensions — formally verified

### v2

@@
  Abstract:  Farkas established that a system of linear inequalities has a solution if and only if we cannot obtain a contradiction by taking a linear combination of the inequalities. We state and formally prove several Farkas-like theorems over linearly ordered fields in Lean 4. Furthermore, we extend duality theory to the case when some coefficients are allowed to take “infinite values”. Keywords:  Farkas lemma, linear programming, extended reals, calculus of inductive constructions
@@ 1 Introduction
In optimization and related fields, we often ask whether a given system of linear inequalities has a solution. Duality theorems answer these naturally arising questions as follows: A system of linear inequalities has a solution if and only if we cannot obtain a contradiction by taking a linear combination of the inequalities. When we formulate duality theorems as “either there exists a solution, or there exists a vector of coefficients that tells us how to derive $0<0$ from given inequalities”, they are usually called theorems of alternatives. Two well-known theorems of alternatives are as follows (Farkas [12, 13]; Minkowski [21]).
Theorem (equalityFarkas ).
@@
Let $I,J$ be finite types. Let $F$ be a linearly ordered field. Let $A$ be a matrix of type $(I\!\times\!J)\rightarrow F$. Let $b$ be a vector of type $I\rightarrow F$. Exactly one of the following exists: nonnegative vector $x:J\rightarrow F$ such that $A\cdot x=b$ vector $y:I\rightarrow F$ such that $A^{T}\!\cdot y\geq 0$ and $b\cdot y<0$
Theorem (inequalityFarkas ).
@@
Let $I,J$ be finite types. Let $F$ be a linearly ordered field. Let $A$ be a matrix of type $(I\!\times\!J)\rightarrow F$. Let $b$ be a vector of type $I\rightarrow F$. Exactly one of the following exists: nonnegative vector $x:J\rightarrow F$ such that $A\cdot x\leq b$ nonnegative vector $y:I\rightarrow F$ such that $A^{T}\!\cdot y\geq 0$ and $b\cdot y<0$
@@
For optimization problems, we obtain a correspondence between the optimal value of a linear program and the optimal value of its dual (originally discussed in the context of zero-sum games by Dantzig and von Neumann [30]; later in Gale, Kuhn, Tucker [14]). The strong duality theorem, which is a cornerstone of linear programming, can be stated as follows.
Theorem (StandardLP.strongDuality ).
@@
Let $I$ and $J$ be finite types. Let $F$ be a linearly ordered field. Let $A$ be a matrix of type $(I\times J)\rightarrow F$. Let $b$ be a vector of type $I\rightarrow F$. Let $c$ be a vector of type $J\rightarrow F$. Then $\min\,\{\,c\cdot x~{}|~{}x\geq 0\operatorname{\;\wedge\;}A\cdot x\leq b\,\}=-\min\,\{\,b\cdot y~{}|~{}y\geq 0\operatorname{\;\wedge\;}(-A^{T})\cdot y\leq c\,\}$ (1) holds if at least one of the systems has a solution (very roughly paraphrased).
@@
Using identity $\min\{c\cdot x~{}|~{}\ldots\}=-\max\{-c\cdot x~{}|~{}\ldots\}$ and replacing $c$ with $-c$, eq. (1) can be equivalently transformed to $\max\,\{\,c\cdot x~{}|~{}x\geq 0\operatorname{\;\wedge\;}A\cdot x\leq b\,\}=\min\,\{\,b\cdot y~{}|~{}y\geq 0\operatorname{\;\wedge\;}A^{T}\!\cdot y\geq c\,\}$ (2) which is probably a more familiar formulation of strong LP duality. Later it will be clear why we chose the “$\min\,$/$\,\min$” rather than the more idiomatic “$\max\,$/$\,\min$” formulation. Also, we do not investigate “asymmetric versions” such as: $\max\,\{\,c\cdot x~{}|~{}x\geq 0\operatorname{\;\wedge\;}A\cdot x=b\,\}=\min\,\{\,b\cdot y~{}|~{}A^{T}\!\cdot y\geq c\,\}$ (3) For many more Farkas-like theorems (beyond the scope of this paper), see for example [29].
@@
  • We formally prove several existing duality results (including the three theorems above) in Lean 4. In fact, we prove a more general version of LABEL:equalityFarkas due to Bartl [4]. Section 1.1 says more about it.
  • We establish (and formally prove in Lean 4) a new generalization of LABEL:inequalityFarkas and LABEL:StandardLP.strongDuality to the case when some of the coefficients are allowed to have infinite values. This scenario can be motivated by discrete optimization problems with “hard” constraints (the “hard” constraints declare what has to be be satisfied, whereas “soft” constraints declare what should be optimized). A common way to concisely write down such problems mathematically is to use infinite coefficients in front of the corresponding terms in the objective. Since infinities are not allowed in traditional LPs, “soft” and “hard” constraints need to be handled differently when formulating LP relaxations of such problems (see e.g. [16]). Our work provides a more direct way to formulate such relaxations. Section 1.2 says more about the extensions.
@@ 1.1 Bartl’s generalization
  • nonnegative vector $x:J\rightarrow R$ such that, for all $w:I\rightarrow R$, we have $\sum_{j:J}\;(A~{}w)_{j}\cdot x_{j}=b~{}w$
  • vector $y:I\rightarrow R$ such that $A~{}y\geq 0$ and $b~{}y<0$
@@
In the next generalization (a similar theorem is in [10]), the partially ordered module $I\rightarrow R$ is replaced by a general $R$-module $W$.
Theorem (almostFarkasBartl ).
@@
  • nonnegative vector $x:J\rightarrow R$ such that, for all $w:W$, we have $\sum_{j:J}\;(A~{}w)_{j}\cdot x_{j}=b~{}w$
  • vector $y:W$ such that $A~{}y\geq 0$ and $b~{}y<0$
@@
  • nonnegative vector family $x:J\rightarrow V$ such that, for all $w:W$, we have $\sum_{j:J}\;(A~{}w)_{j}\cdot x_{j}=b~{}w$
  • vector $y:W$ such that $A~{}y\geq 0$ and $b~{}y<0$
In the last branch, $A~{}y\geq 0$ uses the partial order on $(J\rightarrow R)$ whereas $b~{}y<0$ uses the linear order on $V$. Note that LABEL:fintypeFarkasBartl subsumes LABEL:scalarFarkas (as well as the other versions based on equality), since $R$ can be viewed as a linearly ordered module over itself. We prove LABEL:fintypeFarkasBartl in Section 4, which is where the heavy lifting comes. Our proof is based on [5].
@@ 00     $\bot$ 00 00
[…]
Theorem (extendedFarkas ).
@@
[…]
@@
  • $A$ does not have $\bot$ and $\top$ in the same row
  • $A$ does not have $\bot$ and $\top$ in the same column
  • $A$ does not have $\bot$ in any row where $b$ has $\bot$
  • $A$ does not have $\top$ in any column where $c$ has $\bot$
  • $A$ does not have $\top$ in any row where $b$ has $\top$
  • $A$ does not have $\bot$ in any column where $c$ has $\top$
@@
Theorem (ValidELP.strongDuality ).
@@
$\min\,\{\,c\cdot x~{}|~{}x\geq 0\operatorname{\;\wedge\;}A\cdot x\leq b\,\}=-\min\,\{\,b\cdot y~{}|~{}y\geq 0\operatorname{\;\wedge\;}(-A^{T})\cdot y\leq c\,\}$
If all entries of $(A,b,c)$ are finite then this equation can be stated in many equivalent ways, e.g. as in (2). This reformulation uses the facts that $(-c)\cdot x=-(c\cdot x)$, and that $(-A^{T})\cdot y\leq c$ is equivalent to $A^{T}\cdot y\geq-c$. These facts are no longer true if infinities are allowed, so one must be careful when formulating the duality theorem for $F_{\infty}$. We considered several “$\max\,$/$\,\min$” and “$\max\,$/$\,\max$” versions, but they all required stronger preconditions compared to the “$\min\,$/$\,\min$” version given in LABEL:ExtendedLP.strongDuality.
@@ 1.2.1 Example — cheap lunch
According to [1] a kilogram of boiled white rice contains $27$ g protein and $1300$ kcal. According to [2] a kilogram of boiled lentils contains $90$ g protein and $1150$ kcal. Let’s say that a kilogram of boiled white rice costs $0.92$ euro and that a kilogram of boiled lentils costs $1.75$ euro. We want to cook a lunch, as cheap as possible, that contains at least $30$ g protein and at least $700$ kcal. The choice of white rice and lentils isn’t random — one of the authors ate this lunch at a mathematical camp — at the moment, he didn’t like the lunch — it wasn’t very tasty — but later he realized what an awesome nutritional value the lunch had for its low price. This is a simple example of the well-known diet problem [11].
@@
Using a numerical LP solver, we obtain the optimal solution $p=0.0141594$ and $k=0.000413613$ leading to the objective value $-0.714311$ here. To summarize, the cheapest lunch that contains at least 30 g protein and at least 700 kcal consists of $332$ g white rice and $224$ g lentils; the shadow cost of a gram of protein is $0.014$ euro and the shadow cost of a kcal is $0.00041$ euro in our settings.
Now consider settings where lentils are out of stock. We still want to obtain at least 30 g protein and at least 700 kcal. What is the price of our lunch now? On paper, it is most natural to entirely remove lentils out of the picture and recompute the LP. However, computer proof systems generally don’t like it when sizes of matrices change, so let’s do the modification of input in place. One could suggest that, in order to model the newly arised situation, the price of lentils will be increased to $999999$ euro, so that the optimal solution will assign 00 to it. This is, however, not mathematically elegant, as it requires some engineering insight into setting the constant to an appropriately large number. We claim that the proper mathematical approach is to increase the price of lentils to infinity! Let’s see how our original LP will look now:
@@
At this point, the shadow price of a gram of protein is $0.034074$ euro but the shadow price of kcal is 00 euro. The objective value is $-1.022222$ in accordance with our theoretical prediction.
@@ 1.3 Structure of this paper
In Section 2, we explain all underlying definitions and comment on the formalization process; following the philosophy of [24] we review most of the declarations needed for the reader to believe our results, leaving out many declarations that were used in order to prove the results. In Section 3, we formally state theorems from Section 1 using definitions from Section 2. In Section 4, we prove LABEL:fintypeFarkasBartl (stated in Section 1.1), from which we obtain LABEL:equalityFarkas as a corollary. In Section 5, we prove LABEL:extendedFarkas (stated in Section 1.2). In Section 6, we prove LABEL:ExtendedLP.strongDuality (stated in Section 1.2), from which we obtain LABEL:StandardLP.strongDuality as a corollary. In Section 7, we show what happens when various preconditions are not satisfied.
Repository https://github.com/madvorak/duality/tree/v3.2 contains the full version of all definitions, statements, and proofs. They are written in a formal language called Lean, which provides a guarantee that every step of every proof follows from valid logical axioms.111 The only axioms used in our proofs are propext, Classical.choice, and Quot.sound, which you can check by the #print axioms command. We use Lean version 4.18.0 together with the Lean mathematical library Mathlib [20] revision aa936c3 (dated 2025-04-01). This paper attempts to be an accurate description of the duality project. However, in case of any discrepancy, the code shall prevail.
@@ 2 Preliminaries
There are many layers of definitions that are built before say what a linearly ordered field is (used e.g. in LABEL:equalityFarkas), what a linearly ordered division ring is (used e.g. in LABEL:scalarFarkas), and what a linearly ordered abelian group is (used in LABEL:fintypeFarkasBartl). Section 2.1 mainly documents existing Mathlib definitions (but omitting declarations of default values); we also highlight how we added linearly ordered division rings into our project. Section 2.2 then explains how we defined extended linearly ordered fields. Section 2.3 reviews vectors and matrices; we especially focus on how multiplication between them is defined. Section 2.4 explains how modules are implemented. Section 2.5 provides the formal definition of extended linear programs.

### v3

# Duality theory in linear optimization and its extensions: formally verified

@@
Abstract
Farkas established that a system of linear inequalities has a solution if and only if we cannot obtain a contradiction by taking a linear combination of the inequalities. We state and formally prove several Farkas-like theorems over linearly ordered fields in Lean 4. Furthermore, we extend duality theory to the case when some coefficients are allowed to take “infinite values”.
keywords:
Farkas lemma, linear programming, extended reals, calculus of inductive constructions
[…]
@@ 1 Introduction
In optimization and related fields, we often ask whether a given system of linear inequalities has a solution. Duality theorems answer these naturally arising questions as follows: A system of linear inequalities has a solution if and only if we cannot obtain a contradiction by taking a linear combination of the inequalities. When we formulate duality theorems as “either there exists a solution, or there exists a vector of coefficients that tells us how to derive $0<0$ from given inequalities”, they are usually called theorems of alternatives. Two well-known theorems of alternatives are as follows (Farkas [Farkas1894, Farkas1898]; Minkowski [Minkowski]).
Theorem 1.1 (equalityFarkas ).
@@
Let $I,J$ be finite types. Let $F$ be a linearly ordered field. Let $A$ be a matrix of type $(I\!\times\!J)\rightarrow F$. Let $b$ be a vector of type $I\rightarrow F$. Exactly one of the following exists:
  • nonnegative vector $x:J\rightarrow F$ such that $A\cdot x=b$
  • vector $y:I\rightarrow F$ such that $A^{T}\!\cdot y\geq 0$ and $b\cdot y<0$
Theorem 1.2 (inequalityFarkas ).
@@
Let $I,J$ be finite types. Let $F$ be a linearly ordered field. Let $A$ be a matrix of type $(I\!\times\!J)\rightarrow F$. Let $b$ be a vector of type $I\rightarrow F$. Exactly one of the following exists:
  • nonnegative vector $x:J\rightarrow F$ such that $A\cdot x\leq b$
  • nonnegative vector $y:I\rightarrow F$ such that $A^{T}\!\cdot y\geq 0$ and $b\cdot y<0$
@@
For optimization problems, we obtain a correspondence between the optimal value of a linear program and the optimal value of its dual (originally discussed in the context of zero-sum games by Dantzig and von Neumann [Discussion]; later in Gale, Kuhn, Tucker [GaleKuhnTucker]). The strong duality theorem, which is a cornerstone of linear programming, can be stated as follows.
Theorem 1.3 (StandardLP.strongDuality ).
@@
Let $I$ and $J$ be finite types. Let $F$ be a linearly ordered field. Let $A$ be a matrix of type $(I\times J)\rightarrow F$. Let $b$ be a vector of type $I\rightarrow F$. Let $c$ be a vector of type $J\rightarrow F$. Then
$\min\,\{\,c\cdot x\penalty 10000\ |\penalty 10000\ x\geq 0\operatorname{\;\wedge\;}A\cdot x\leq b\,\}=-\min\,\{\,b\cdot y\penalty 10000\ |\penalty 10000\ y\geq 0\operatorname{\;\wedge\;}(-A^{T})\cdot y\leq c\,\}$ (1)
holds if at least one of the systems has a solution (very roughly paraphrased).
Using identity $\min\{c\cdot x\penalty 10000\ |\penalty 10000\ \ldots\}=-\max\{-c\cdot x\penalty 10000\ |\penalty 10000\ \ldots\}$ and replacing $c$ with c-c, eq. (1) can be equivalently transformed to
$\max\,\{\,c\cdot x\penalty 10000\ |\penalty 10000\ x\geq 0\operatorname{\;\wedge\;}A\cdot x\leq b\,\}=\min\,\{\,b\cdot y\penalty 10000\ |\penalty 10000\ y\geq 0\operatorname{\;\wedge\;}A^{T}\!\cdot y\geq c\,\}$ (2)
@@
which is probably a more familiar formulation of strong LP duality. Later it will be clear why we chose the $\min\,$$\min$” rather than the more idiomatic $\max\,$$\min$” formulation. Also, we do not investigate “asymmetric versions” such as:
$\max\,\{\,c\cdot x\penalty 10000\ |\penalty 10000\ x\geq 0\operatorname{\;\wedge\;}A\cdot x=b\,\}=\min\,\{\,b\cdot y\penalty 10000\ |\penalty 10000\ A^{T}\!\cdot y\geq c\,\}$ (3)
For many more Farkas-like theorems (beyond the scope of this paper), see for example [Perng2017].
@@
  • We formally prove several existing duality results (including the three theorems above) in Lean 4. In fact, we prove a more general version of LABEL:equalityFarkas due to Bartl [Bartl2007]. Section 1.1 says more about it.
  • We establish (and formally prove in Lean 4) a new generalization of LABEL:inequalityFarkas and LABEL:StandardLP.strongDuality to the case when some of the coefficients are allowed to have infinite values. This scenario can be motivated by discrete optimization problems with “hard” constraints (the “hard” constraints declare what has to be be satisfied, whereas “soft” constraints declare what should be optimized). A common way to concisely write down such problems mathematically is to use infinite coefficients in front of the corresponding terms in the objective. Since infinities are not allowed in traditional LPs, “soft” and “hard” constraints need to be handled differently when formulating LP relaxations of such problems (see e.g. [BLP]). Our work provides a more direct way to formulate such relaxations. Section 1.2 says more about the extensions.
@@ 1.1 Bartl’s generalization
  • nonnegative vector $x:J\rightarrow R$ such that, for all $w:I\rightarrow R$, we have $\sum_{j:J}\;(A\penalty 10000\ w)_{j}\cdot x_{j}=b\penalty 10000\ w$
  • vector $y:I\rightarrow R$ such that $A\penalty 10000\ y\geq 0$ and $b\penalty 10000\ y<0$
@@
In the next generalization (a similar theorem is in [Chernikov]), the partially ordered module $I\rightarrow R$ is replaced by a general $R$-module $W$.
Theorem 1.5 (almostFarkasBartl ).
@@
  • nonnegative vector $x:J\rightarrow R$ such that, for all $w:W$, we have $\sum_{j:J}\;(A\penalty 10000\ w)_{j}\cdot x_{j}=b\penalty 10000\ w$
  • vector $y:W$ such that $A\penalty 10000\ y\geq 0$ and $b\penalty 10000\ y<0$
@@
  • nonnegative vector family $x:J\rightarrow V$ such that, for all $w:W$, we have $\sum_{j:J}\;(A\penalty 10000\ w)_{j}\cdot x_{j}=b\penalty 10000\ w$
  • vector $y:W$ such that $A\penalty 10000\ y\geq 0$ and $b\penalty 10000\ y<0$
In the last branch, $A\penalty 10000\ y\geq 0$ uses the partial order on $(J\rightarrow R)$ whereas $b\penalty 10000\ y<0$ uses the linear order on $V$. Note that LABEL:fintypeFarkasBartl subsumes LABEL:scalarFarkas (as well as the other versions based on equality), since $R$ can be viewed as a linearly ordered module over itself. We prove LABEL:fintypeFarkasBartl in Section 4, which is where the heavy lifting comes. Our proof is based on [Bartl2011].
@@ 0     $\bot$ 0 0
Informally speaking, $\top$ represents the positive infinity, $\bot$ represents the negative infinity, and we say that $\bot$ is “stronger” than $\top$ in all arithmetic operations. The surprising parts are $\bot+\top=\bot$ and $0\thinspace\cdot\bot=\bot$. Because of them, $F_{\infty}$ is not a field. In fact, $F_{\infty}$ is not even a group. However, $F_{\infty}$ is still a densely linearly ordered abelian monoid with characteristic zero. Note that $\bot+\top=\bot$ is a standard convention in Mathlib [Mathlib] but $0\thinspace\cdot\bot=\bot$ is ad hoc.
Theorem 1.8 (extendedFarkas ).
@@
Note LABEL:extendedFarkas has four preconditions on matrix $A$ and vector $b$. In Section 7.1 we show that omitting any of them makes the theorem false. Observe that LABEL:inequalityFarkas has condition $A^{T}\cdot y\geq 0$ in the second branch, while in LABEL:extendedFarkas we changed it to $(-A^{T})\cdot y\leq 0$. The two conditions are equivalent for finite-valued matrices $A$, but not necessarily for matrices $A$ with infinities (e.g. condition $(-\top)\cdot 0\geq 0$ is false but $\top\cdot 0\leq 0$ is true). One must be careful when formulating this condition; for example, using the condition from LABEL:inequalityFarkas would make the theorem false even if $A$ has only a single $\bot$ entry (see Section 7.1).
@@
  • AA does not have $\bot$ and $\top$ in the same row
  • AA does not have $\bot$ and $\top$ in the same column
  • AA does not have $\bot$ in any row where $b$ has $\bot$
  • AA does not have $\top$ in any column where $c$ has $\bot$
  • AA does not have $\top$ in any row where $b$ has $\top$
  • AA does not have $\bot$ in any column where $c$ has $\top$
@@
Theorem 1.11 (ValidELP.strongDuality ).
@@
$\min\,\{\,c\cdot x\penalty 10000\ |\penalty 10000\ x\geq 0\operatorname{\;\wedge\;}A\cdot x\leq b\,\}=-\min\,\{\,b\cdot y\penalty 10000\ |\penalty 10000\ y\geq 0\operatorname{\;\wedge\;}(-A^{T})\cdot y\leq c\,\}$
If all entries of $(A,b,c)$ are finite then this equation can be stated in many equivalent ways, e.g. as in (2). This reformulation uses the facts that $(-c)\cdot x=-(c\cdot x)$, and that $(-A^{T})\cdot y\leq c$ is equivalent to $A^{T}\cdot y\geq-c$. These facts are no longer true if infinities are allowed, so one must be careful when formulating the duality theorem for $F_{\infty}$. We considered several $\max\,$$\min$” and $\max\,$$\max$” versions, but they all required stronger preconditions compared to the $\min\,$$\min$” version given in LABEL:ExtendedLP.strongDuality.
@@ 1.2.1 Example—cheap lunch
According to Nutritionix [^1] https://www.nutritionix.com/food/white-rice/1000-g a kilogram of boiled white rice contains $27$ g protein and $1300$ kcal. According to Nutritionix [^2] https://www.nutritionix.com/food/lentils a kilogram of boiled lentils contains $90$ g protein and $1150$ kcal. Let’s say that a kilogram of boiled white rice costs $0.92$ euro and that a kilogram of boiled lentils costs $1.75$ euro. We want to cook a lunch, as cheap as possible, that contains at least $30$ g protein and at least $700$ kcal. The choice of white rice and lentils isn’t random — one of the authors ate this lunch at a mathematical camp — at the moment, he didn’t like the lunch — it wasn’t very tasty — but later he realized what an awesome nutritional value the lunch had for its low price. This is a simple example of the well-known diet problem [Dantzig].
@@
Using a numerical LP solver, we obtain the optimal solution $p=0.0141594$ and $k=0.000413613$ leading to the objective value 0.714311-0.714311 here. To summarize, the cheapest lunch that contains at least 30 g protein and at least 700 kcal consists of $332$ g white rice and $234$ g lentils; the shadow cost of a gram of protein is $0.014$ euro and the shadow cost of a kcal is $0.00041$ euro in our settings.
Now consider settings where lentils are out of stock. We still want to obtain at least 30 g protein and at least 700 kcal. What is the price of our lunch now? On paper, it is most natural to entirely remove lentils out of the picture and recompute the LP. However, computer proof systems generally don’t like it when sizes of matrices change, so let’s do the modification of input in place. One could suggest that, in order to model the newly arised situation, the price of lentils will be increased to $999999$ euro, so that the optimal solution will assign 0 to it. This is, however, not mathematically elegant, as it requires some engineering insight into setting the constant to an appropriately large number. We claim that the proper mathematical approach is to increase the price of lentils to infinity! Let’s see how our original LP will look now:
@@
At this point, the shadow price of a gram of protein is $0.034074$ euro but the shadow price of kcal is 0 euro. The objective value is 1.022222-1.022222 in accordance with our theoretical prediction.
@@ 1.3 Structure of this paper
In Section 2, we explain all underlying definitions and comment on the formalization process; following the philosophy of [Believe] we review most of the declarations needed for the reader to believe our results, leaving out many declarations that were used in order to prove the results. In Section 3, we formally state theorems from Section 1 using definitions from Section 2. In Section 4, we prove LABEL:fintypeFarkasBartl (stated in Section 1.1), from which we obtain LABEL:equalityFarkas as a corollary. In Section 5, we prove LABEL:extendedFarkas (stated in Section 1.2). In Section 6, we prove LABEL:ExtendedLP.strongDuality (stated in Section 1.2), from which we obtain LABEL:StandardLP.strongDuality as a corollary. In Section 7, we show what happens when various preconditions are not satisfied.
Repository https://github.com/madvorak/duality/tree/v3.2 contains the full version of all definitions, statements, and proofs. They are written in a formal language called Lean, which provides a guarantee that every step of every proof follows from valid logical axioms. [^3] The only axioms used in our proofs are propext, Classical.choice, and Quot.sound, which you can check by the #print axioms command. We use Lean version 4.18.0 together with the Lean mathematical library Mathlib [Mathlib] revision aa936c3 (dated 2025-04-01). This paper attempts to be an accurate description of the duality project. However, in case of any discrepancy, the code shall prevail.
@@ 2 Preliminaries
There are many layers of definitions that are built before say what a linearly ordered field is, what a linearly ordered division ring is, and what a linearly ordered abelian group (used in LABEL:fintypeFarkasBartl) is. Section 2.1 mainly documents existing Mathlib definitions (but omitting declarations of default values); we also highlight how we added linearly ordered division rings into our project. Section 2.2 then explains how we defined extended linearly ordered fields. Section 2.3 reviews vectors and matrices; we especially focus on how multiplication between them is defined. Section 2.4 explains how modules are implemented. Section 2.5 provides the formal definition of extended linear programs.